# Strategic resource allocation in memory encoding: An efficiency principle shaping language processing

### v1

@@ 5.2 Results
Figure 7: Study 2b reading time (RT) result for first-pass duration and total RT. Panel A: raw RTs at the right codependent as a function of dependency length $L$ modulated by antecedent surprisal, which is binned into tertiles for visualization. Panel B: result of regression models with log-transformed RTs; coefficient estimates with 95% confidence interval for the effects of $L$, antecedent surprisal, and the interaction between the two. Significance levels: *(p$<$0.05), **(p$<$0.01), ***(p$<$0.001)
Figure 7A shows the interaction effect between dependency length $L$ and antecedent surprisal on raw reading times, including both the first-pass duration and the total RT. The result of statistical models for the critical effects is summarized in Figure 7B.
@@
As shown in Figure 7B (left column), no evidence was found for the main effect of dependency length $L$ with any measure of RT and $L$, suggesting the lack of the baseline locality effect. No significant main effect of antecedent surprisal, either. For the critical $L$$\times$ antecedent surprisal interaction, although the effect is numerically negative on first-pass RT with both measures of $L$, it is only marginally significant. No evidence for the interaction effect was found on total RT.
@@ 5.3 Discussion
The result of the two main effects in the current Study 2b shows a very different pattern compared to Study 2a. First, unlike Study 2a, in the current Study 2b only in object relations did we find a baseline locality effect, whereas the locality effect is not observed in the analysis of the full dataset or in the one of subject relations. That is, longer dependency length does not lead to higher processing difficulty at the retrieval site for subject relations. This lack of the baseline locality effect aligns with the observation in Demberg and Keller (2008), where the locality effect in Dundee corpus is overall small and unreliable for verbs, which in our case is the retrieval site of subject relations. Second, there is no main effect of antecedent surprisal across the board in this Study 2b, casting doubt on how reliable the negative antecedent surprisal main effect is in Study 2a.
@@ 6 General Discussion
In three corpus studies, we examined strategic resource allocation through the lens of dependency locality both in production and in comprehension. Study 1 explored this hypothesis in production by analyzing UD corpora from 10 languages. Our result reveals that more surprising antecedents can tolerate more intervening material before they need to be retrieved at the other side of the dependency, resulting in a positive correlation between antecedent surprisal and dependency length. However, it is worth noting that this reduced locality effect mostly exists within Indo-European and head-initial languages, and is more consistent for dependencies of subject relations than for object relations.
@@
Overall, despite the caveats mentioned above, our result shows converging evidence that a reduced locality effect emerges for more surprising antecedents in the processing of non-local dependencies, suggesting that more surprising antecedents are less susceptible to the interference from intervening material. This finding aligns with the notion of strategic resource allocation that we proposed, which holds that unexpected information is prioritized for memory resources and is encoded with enhanced memory representation.
@@ 6.1 The role of context in working memory efficiency
In the current study, that the observed antecedent surprisal effect cannot be reduced to a pure frequency effect suggests that the statistics relevant to working memory efficiency go beyond unigram frequencies, and the strategic allocation of working memory resources is based on more fine-grained context-specific probabilities. As noted by Jaeger & Buz (2017), frequency can be understood as an averaged effect of contextual predictability. This is probably one of the reasons why linguistic theories focusing on the representational aspects of language often emphasize frequency, as it reflects the abstract, global properties of a language accumulated through long-term experience. That being said, the optimization of working memory efficiency not only relies on those ready-to-retrieve statistics already stored in long-term memory, but also incorporates statistics computed online, dynamically drawing upon rich contextual information in a rapid and adaptive manner (for a different view, see Opedal et al., 2024). It is worth noting that by using surprisals from autoregressive language models the current study only tests the left context of a linguistic unit. A natural direction for future work is to explore the role of bidirectional context in working memory efficiency.
@@ 6.2 Relationship with lossy-context surprisal
The mechanism of strategic resource allocation we proposed bears similar theoretical and empirical implications with lossy-context surprisal theory and its variants (Futrell, Gibson, and Levy, 2020; Hahn et al., 2022). Focusing on the prediction mechanism, lossy-context surprisal holds that next-word predictions are based on lossy and faulty memory representations, rather than the veridical representation of the past linguistic input. The theory explicitly includes a memory distortion process, where certain elements in an utterance are erased to form a lossy representation, subject to certain probability distributions.
@@
In many aspects, our proposal and lossy-context surprisal form two sides of the same coin. For us, we seek to understand and explain the working memory mechanism in sentence processing, and the mechanism we proposed includes a component of predictive processing. For lossy-context surprisal, the theory aims to understand the prediction mechanism in sentence processing, which includes a memory component. However, we both speak for an interaction between memory and prediction, as both of them should jointly support human linguistic behaviors as a cognitive task. That means, our proposal and lossy-context surprisal are not mutually exclusive, theoretically or empirically, and we simply focus on different perspectives of a similar cognitive task.
@@
In fact, our theory of strategic resource allocation and lossy-context surprisal are on some level mutually translatable, so we do not see them necessarily as conflicting theories. On the one hand, in the language of lossy-context surprisal, the memory distortion process where certain linguistic units are erased is potentially where our strategic resource allocation could fit in, such that more surprising units given the context are more likely to be erased when predicting future units. On the other hand, the memory component in lossy-context surprisal also fits in the resource-rational explanation of working memory efficiency. As outlined above in Introduction, a fundamental question at the core of our resource-rational analysis of working memory mechanism is: if memory capacity is limited, how to minimize the distortion cost by strategically prioritizing more important linguistic units? However, what counts as important? Or, in other words, what is the objective function based on which the distortion cost is defined? In the original lossy-context surprisal theory (Futrell, Gibson, and Levy, 2020), this objective function is not explicitly specified. In Hahn et al. (2022), the model takes one step further, and this objective function is to minimize the downstream next-word prediction task. In our proposal, the distortion cost is defined by how likely a unit can be reconstructed later given the context. These two different objective functions of distortion cost are not mutually exclusive, and the accuracy of next-word prediction may be compromised if a lost unit in memory is not reconstructable. Therefore, although our proposal of strategic resource allocation has a different theoretical focus from the model of Hahn et al. (2022), our empirical predictions actually share some overlap, and we both predict on some level that the representation of more surprising units (or, less frequeny units) should be more robust (or, less likely to be deleted in the lossy-context surprisal model).
@@ 6.3 Hierarchical encoding and compression
Strategic resource allocation, arising as an efficiency principle from the functional pressures of working memory, is situated more at the computational level in Marr (1982)’s three-level representation. A natural question to ask then is: what is the potential mechanism to implement this efficiency principle at the algorithmic level?
@@
Predictable information, if less prioritized, should be encoded in a more compressed and abstract fashion. In visual working memory, this is indeed evidenced by the fact that more items in a visual memory task can be stored when their perceptual features are statistically correlated (Brady et al., 2009; Bates et al., 2019). Moreover, information with stronger prior knowledge is also more susceptible to the categorical bias in perception, where the perceived input is biased towards the categorical mean (Bates and Jacobs, 2020). In the domain of language, more frequent linguistic sequence is more likely to be holistically stored in memory (e.g., Bybee, 2006; Goldberg, 2003; Hawkins, 2004; Traugott & Trousdale, 2013). Importantly, hierarchical compression based on statistical regularities provides a mechanism that strings together effects across different linguistic representational levels, forming a spectrum of compression. At one end, there is the locality effect where words that are more mutually predictable tend to stay closer to each other in linear order (e.g., Futrell et al., 2019; Futrell, 2019). At the other end, the same pressure of compression governs the fusion of morphemes (e.g., Hahn et al., 2021; Rathi et al., 2021), and makes mutually predictable units more likely to go through processes such as affixation and phonological reduction (e.g., Bybee et al., 1994; Bybee, 2006; Gahl and Baayen, 2024).
@@ 6.4 Parallelism between production and comprehension
The effect of strategic resource allocation holds both for production and for comprehension in the current study, pointing to the parallelism between these two modalities. For comprehension, a reasonable question is: to what extent the observed effect of strategic resource allocation is experience-based (MacDonald & Christiansen, 2002), given that the same effect is also seen in production data? In other words, it is possible that comprehenders prioritize unexpected linguistic units and encode them with enhanced representation because unexpected units are more likely to associate with stronger memory interference in the production data they receive.
A similar question can also be asked for production: to what extent the effect observed in production is the result of audience design (H. H. Clark and Murphy, 1982; Ferreira, 2019; Lockridge & Brennan, 2002), given that comprehenders can encode unexpected units with prioritized memory resources. One listner-oriented production theory compatible with our finding is the Uniform Information Density (UID) theory (Jaeger & Levy, 2006; Meister et al., 2021; T. H. Clark et al., 2023). According to UID, surprising antecedents may be followed by longer dependency length so that there is a smoother transition to the other side of the dependency, which is relatively predictable since it shares high mutual information with the antecedent (Futrell et al., 2019).
Here we do not attempt to adjudicate the possibilities mentioned above, nor do we view these possibilities in conflict with our proposal. In our opinion, strategic resource allocation provides a potential explanation for the mechanistic underpinnings of these higher-level processes.
Despite the parallelism, comprehension and production still differ in some critical aspects, exerting modality-specific constraints on strategic resource allocation due to their idiosyncratic processing nature. For example, production is, in general,l, more cognitively demanding, in need of higher memory capacity, executive control, and action planning than comprehension (MacDonald, 2013; Nozari & Novick, 2017; Hickok, 2012; Koranda et al., 2020). This additional cognitive demand may exert more pressure to efficiently use the limited memory resources in production than in comprehension, possibly resulting in a stronger effect of strategic resource allocation. Future work is needed to investigate this possibility.
@@ 6.5 Implications for linguistic typology
Dependency length minimization as a functional universal has been argued to shape the syntactic structure of human language. As reviewed in Introduction, this universal arises from the pressure to efficiently use the limited working memory resources. In the current study, we go beyond the general constraint of limited memory capacity, and further argue that memory resources should be strategically allocated to prioritize novel and unexpected information, a memory efficiency principle that naturally arises from two assumptions about working memory. The results indicate that this strategic resource allocation indeed serves as a functional constraint to shape syntactic structures: although there is a general pressure to minimize dependency length, this pressure can actually be relaxed when the antecedent of a syntactic dependency is of higher surprisal. Our finding further substantiates the functionalist view as a promising approach to provide explanatory accounts for linguistic universals (Gibson et al., 2019). It also highlights the importance of having increasingly sophisticated characterization of functional constraints, in order to see how far this functional perspective of language structure can go.
@@
How does strategic resource allocation as a universal efficiency principle interact with other grammatical constraints and language-specific phrase structures? One consistent pattern we observed is the asymmetry between subject and object relations. Specifically, the effect of strategic resource allocation is generally smaller and less reliable for object relations than for subject relations. One possible explanation is that object relations are subject to stronger grammatical constraints, with greater pressure to position the head and its dependent closer to each other. In support of this possibility, Keenan & Comrie (1977) identify an Accessibility Hierarchy as a linguistic universal, where noun phrases in the subject position are more readily relativized into relative clauses than those in the object position. Such constraints may bind object noun phrases and verbs more tightly, reducing the influence of strategic resource allocation. Moreover, in many languages, grammatical agreement is common in subject-verb dependencies but absent in object-verb dependencies. Therefore, establishing subject relations may require more active grammatical computation, resulting in increased memory demand in processing subject relations and a stronger need for more strategic and efficient use of working memory resources.
@@
Another notable pattern in Study 1 is that, compared to head-initial languages, most of the head-final languages in our analysis do not exhibit a reliable antecedent surprisal effect on dependency length. We speculate that this may be related to the tendency for argument dropping in head-final languages with SOV word order. From the perspective of dependency locality, SOV word order is associated with longer dependency lengths compared to SVO, which should theoretically impose higher memory costs and make it a less efficient structure. Despite this seemingly inefficiency, typologically, SOV is a word order commonly attested (Hammarström, 2016). To address this paradox, some studies find that arguments in SOV languages are often dropped, thus reducing the overall dependency length in actual language use (Levshina, 2025; Ueno & Polinsky, 2009). It is possible that by allowing speakers to drop arguments, the efficiency of SOV structure may already be significantly improved, obviating the need for strategic resource allocation as another efficiency strategy.
@@
6.6 Surprisal effect as efficient coding

### v2

@@ 6.2 Results
Figure 10: Study 2b reading time (RT) result for first-pass duration and total RT. Panel A: raw RTs at the right codependent as a function of dependency length $L$ modulated by antecedent surprisal, which is binned into tertiles for visualization. Panel B: result of regression models with log-transformed RTs; coefficient estimates with 95% confidence interval for the effects of $L$, antecedent surprisal, and the interaction between the two. Significance levels: *(p<<0.05), **(p<<0.01), ***(p<<0.001)
Figure 10A shows the interaction effect between dependency length $L$ and antecedent surprisal on raw reading times, including both the first-pass duration and the total RT. The result of statistical models for the critical effects is summarized in Figure 10B. [^17] No critical effects were found in the spillover region, so we only report the result of the critical region in this Study 2b.
@@
As shown in Figure 10B (left column), no evidence was found for the main effect of dependency length $L$ with any measure of RT and $L$, suggesting the lack of the baseline locality effect. No significant main effect of antecedent surprisal was found, either. For the critical $L$ $\times$ antecedent surprisal interaction, although the effect is numerically negative on first-pass RT with both measures of $L$, it is only marginally significant. No evidence for the interaction effect was found on total RT.
@@ 6.3 Discussion
The result of the two main effects in the current Study 2b shows a very different pattern compared to Study 2a. First, unlike Study 2a, in the current Study 2b only in object relations did we find a baseline locality effect, whereas the locality effect is not observed in the analysis of the full dataset or in the one of subject relations. That is, longer dependency length does not lead to higher processing difficulty at the retrieval site for subject relations. This lack of the baseline locality effect aligns with the observation in Demberg and Keller (2008), where the locality effect in Dundee corpus is overall small and unreliable for verbs, which in our case is the retrieval site of subject relations. Second, there is no main effect of antecedent surprisal across the board in this Study 2b.
@@ 7 General Discussion
In three corpus studies, we examined strategic resource allocation (SRA) through the lens of dependency locality both in production and in comprehension. Study 1 explored this hypothesis in production by analyzing UD corpora of 10 languages. Our result reveals that more surprising antecedents can tolerate more intervening material before they need to be retrieved at the other side of the dependency, resulting in a positive correlation between antecedent surprisal and dependency length. However, it is worth noting that this reduced locality effect mostly exists within Indo-European and head-initial languages, and is more consistent for dependencies of subject relations than for object relations.
@@
Overall, despite the caveats mentioned above, our result shows emerging evidence that a reduced locality effect emerges for more surprising antecedents in the processing of non-local dependencies, suggesting that more surprising antecedents are less susceptible to the interference from intervening material. This finding aligns with the notion of strategic resource allocation that we proposed, which holds that unexpected information is prioritized for memory resources and is encoded with enhanced memory representation.
7.1 Processing difficulty as encoding difficulty: Reinterpreting the effect of surprisal and entropy
In this section, we first reinterpret the processing difficulty of a word as its memory encoding difficulty. As shown in Figure 11, the encoding process in Figure 2 can be considered a transformation from a flat uniform distribution over all possible words to one that is concentrated around the received input. The processing difficulty of a word at the encoding stage, therefore, can be considered the distance between the pre-encoding and the post-encoding distribution. As a result, less uncertain, or more precise encoding distribution is more distant away from the uniform pre-encoding distribution, leading to higher processing difficulty. With this encoding view in mind, let us consider two factors that have been previously argued to influence the processing difficulty of a word, namely surprisal and entropy.
Figure 11: Processing difficulty as memory encoding difficulty
For the surprisal effect, as shown in our main proposal of SRA, more surprising input should be encoded with higher precision, an efficiency strategy that we have argued to minimize the retrieval error at a later time point. As a result, SRA naturally predicts that the more precise encoding for more surprising inputs should lead to higher encoding or processing difficulty, which is consistent with the widely observed surprisal effect in the literature. Importantly, this memory encoding view provides a resource-rational account for the surprisal effect, reinterpreting it as a strategic solution to the efficiency problem of memory. We will discuss this in more detail below in Section 7.8.
For the effect of entropy, SRA yields complicated predictions. As mentioned in Introduction, SRA implies that higher uncertainty of prior does not necessarily increase or decrease the precision of encoding distribution, and therefore does not necessarily increase or decrease the processing difficulty of a word. To demonstrate the reason behind it, first recall that according to SRA the encoding precision depends on how the received input can be accurately reconstructed at a later time point, which is in turn dependent on to what extent such a reconstruction can be supported by the prior. Intuitively, when the prior is highly consistent with the actual received word (e.g., low surprisal words), the more precise the prior is, the less precise the encoding of that word needs to be for better reconstruction. In this case, the uncertainty of prior should have a positive effect on encoding difficulty, in the sense that the processing of a word is facilitated if its preceding context yields a high-constraining prior. However, when the prior is not compatible with the actual input (e.g., high surprisal words), such a facilitation effect is necessarily the case any more: a highly precise but incompatible prior may actually need to be counteracted by more effort into precisely encode the actual input. In fact, this complicated effect of prior precision echos the empirical observation in many previous studies, where the effect of prediction entropy is much less reliable than the effect of surprisal on processing difficulty (Linzen & Jaeger, 2016; van Schijndel & Linzen, 2021; Wilcox et al., 2023).
7.2 The role of representational uncertainty
Under SRA, there is a dissociation between accuracy and uncertainty in memory representations. For accuracy, it is more related to the mean, or point estimate of an underlying distribution with respect to how far it is from the true input value. For uncertainty, it corresponds to the precision, or variance of the distribution, reflecting the relative competitiveness of all alternative inputs. As mentioned in Introduction, the critical prediction of SRA is about the precision or uncertainty in the encoded memory representation, as reflected in robustness to interference, rather than about the raw accuracy of retrieval. As shown in Figure 2, although a relatively high retrieval accuracy may be maintained on average for all linguistic inputs, those that receive more resources will have less uncertainty in their encoded representation.
Both the point estimate and the uncertainty are important information to understand the underlying representations of memory processes, as raised by more and more recent work in psychophysics (Bays et al., 2024). However, compared to point estimates, the characterization of representational uncertainty is relatively underexplored in psycholinguistics research, both theoretically and empirically. For example, studies under the framework of cue-based retrieval often concern what representation has actually been retrieved, based on which an empirical prediction is derived. Similarly, studies under the noisy-channel framework often focus more on interpreting the point estimate of posterior distribution, rather than how the probability mass is distributed over the hypothesis space.
One of the major challenges to understand the role of representational uncertainty is possibly the lack of a straightforward linking hypothesis. In most psycholinguistics studies, the interpretation of online dependent measures such as reading times is based on the point estimate of its mean, which is naturally linked to the point estimate of mental representations (cf. Huang & Dillon, 2023). But for representational uncertainty, how this psychological construct is linked to any online behavioral measure remains unclear. Even though there might be some ways to probe the degree of representational uncertainty through certain offline measures (e.g., tasks that directly probe the errors in the interpretation of a sentence), it is still challenging to lay out the hypothesis space of alternative representations in a fine-grained manner.
In the current study, an important assumption we made is that less uncertainty leads to more robust representation against interference. That is, if the processesor is more uncertain about the representation of a linguistic input, its encoding distribution may be more likely to be influenced and distorted by other information it holds in memory. Surely, this is a debatable assumption, and the specific mechanism of how the representation is distorted by other inputs needs to be further elaborated in future work.
@@ 7.3 The role of context in working memory efficiency
In the current study, that the observed antecedent surprisal effect cannot be reduced to a pure frequency effect suggests that the statistics relevant to working memory efficiency go beyond unigram frequencies, and the strategic allocation of working memory resources is based on more fine-grained context-specific probabilities. As noted by Jaeger & Buz (2017), frequency can be understood as an averaged effect of contextual predictability. This is probably one of the reasons why linguistic theories focusing on the representational aspects of language often emphasize frequency, as it reflects the abstract, global properties of a language accumulated through long-term experience. That being said, the optimization of working memory efficiency not only relies on those ready-to-retrieve statistics already stored in long-term memory, but also incorporates statistics computed online, dynamically drawing upon rich contextual information in a rapid and adaptive manner (for a different view, see Opedal et al., 2024).
Moreover, the role of context in SRA has its implications for the asymmetry between proactive and retroactive interference. Proactive interference refers to the configuration where distractor information precedes the retrieval target, whereas for retroactive interference the distractor is located between the retrieval target and the retrieval site. Previous studies have observed that proactive interference has weaker effect than the retroactive one (Van Dyke & McElree, 2011), an empirical pattern that supports time-based decay of memory activation (Portrat et al., 2008; Barrouillet et al., 2004; Page & Norris, 1998; Lewis & Vasishth, 2005). Under SRA, this asymmetry can be potentially explained from a resource-rational perspective. As shown in Figure 2, at the encoding stage of a linguistic unit, the processor already has access to the information in its preceding context, but not yet to the upcoming information in the right context. Therefore, during memory encoding, the strategic allocation is possibly based only on the preceding context. In other words, the encoded representation of an input is optimized for information it has already received in the preceding context, but not necessarily for what has not yet been received. If this is the case, SRA naturally explains why the distractor that goes before the target unit has less impact on its representation than the retroactive distractor, without necessarily resorting to a separate time-based decay mechanism.
@@ 7.4 Relationship with lossy-context surprisal
The theoretical framework of SRA we proposed shares similar theoretical and empirical implications with lossy-context surprisal theory and its variants (Futrell, Gibson, and Levy, 2020; Hahn et al., 2022). Focusing on the prediction mechanism, lossy-context surprisal holds that next-word predictions are based on lossy and faulty memory representations, rather than the veridical form of the past linguistic input. The theory explicitly includes a memory distortion process, where certain elements in an utterance are erased to form a lossy representation, subject to certain probabilistic erasure distributions.
@@
In many aspects, our proposal and lossy-context surprisal form two sides of the same coin. For our SRA, we seek to understand and explain the working memory mechanism in sentence processing, and predictive processing is a component incorporated into the mechanism we proposed to better explain memory. For lossy-context surprisal, in contrast, the theory aims to understand the prediction mechanism in sentence processing, with a memory component included to better explain prediction. Despite these different goals, we both speak for an interaction between memory and prediction, as both of them should jointly support human linguistic behaviors as a cognitive task. That means, our proposal and lossy-context surprisal are not mutually exclusive, theoretically or empirically, and we simply focus on different perspectives of a similar cognitive task.
@@
In fact, our theoretical framework of SRA is on some level mutually translatable with lossy-context surprisal, so we do not see them necessarily as conflicting theories. On the one hand, in the language of lossy-context surprisal, the memory distortion process where certain linguistic units are erased is potentially where our strategic resource allocation could fit in, such that more surprising units given the context are less likely to be erased when predicting future units. On the other hand, a fundamental question at the core of our resource-rational analysis of working memory mechanism is: if memory capacity is limited, how to minimize the cost of memory error by strategically prioritizing more important linguistic units? However, what counts as important? Or, in other words, what is the objective function based on which the cost of memory error is defined? In the original lossy-context surprisal theory (Futrell, Gibson, and Levy, 2020), this objective function is not explicitly specified. In Hahn et al. (2022), the model takes one step further, and this objective function is to minimize the downstream next-word prediction task. In our proposal, the cost is defined by how likely a unit can be reconstructed later given the context. These two different objective functions of cost are not mutually exclusive, and the accuracy of next-word prediction may be compromised if a lost unit in memory is not reconstructable. Therefore, although our proposal of SRA has a different theoretical focus from the model of Hahn et al. (2022), our empirical predictions actually share some overlap, and we both predict on some level that the representation of more surprising units (or, less frequent units) should be enhanced and more robust.
@@ 7.5 Hierarchical encoding and compression
SRA, arising as an efficiency principle from the functional pressures of working memory, is situated more at the computational level in Marr (1982)’s three-level representation. A natural question to ask then is: what is the potential mechanism to implement this efficiency principle at the algorithmic level? In other words, when more resources are allocated, what makes the representation less uncertain and more robust?
@@
Predictable information, if less prioritized, should be encoded in a more compressed and abstract fashion. In visual working memory, this is indeed evidenced by the fact that more memory objects can be stored in visual working memory tasks when their perceptual features are statistically correlated (Brady et al., 2009; Bates et al., 2019). Moreover, information with stronger prior knowledge is also more susceptible to the categorical bias in perception, where the perceived input is biased towards the categorical mean (Bates and Jacobs, 2020). Similarly, in the domain of language, more frequent linguistic sequence is more likely to be holistically stored in memory (e.g., Bybee, 2006; Goldberg, 2003; Hawkins, 2004; Traugott & Trousdale, 2013). Importantly, hierarchical compression based on statistical regularities provides a mechanism that strings together effects across different linguistic representational levels, forming a spectrum of compression. At one end, there is the locality effect where words that are more mutually predictable tend to stay closer to each other in linear order (e.g., Futrell et al., 2019; Futrell, 2019). At the other end, the same pressure of compression governs the fusion of morphemes (e.g., Hahn et al., 2021; Rathi et al., 2021), and makes mutually predictable units more likely to go through processes such as affixation and phonological reduction (e.g., Bybee et al., 1994; Bybee, 2006; Gahl and Baayen, 2024).
@@ 7.6 Parallelism between production and comprehension
The effect of SRA holds both for production and for comprehension in the current study, pointing to a parallelism between these two modalities. For comprehension, a reasonable question is: to what extent the observed effect of strategic resource allocation is experience-based (MacDonald & Christiansen, 2002), given that the same effect is also seen in production data? In other words, it is possible that comprehenders prioritize unexpected linguistic units and encode them with enhanced representation because unexpected units are more likely to associate with stronger memory interference in the production data they receive.
A similar question can be asked for production as well: to what extent is the effect observed in production the result of audience design (H. H. Clark and Murphy, 1982; V. S. Ferreira, 2019; Lockridge & Brennan, 2002), given that comprehenders can encode unexpected units with prioritized memory resources? One listener-oriented production theory compatible with our finding is the Uniform Information Density (UID) theory (Jaeger & Levy, 2006; Meister et al., 2021; T. H. Clark et al., 2023). According to UID, surprising antecedents may be followed by longer dependency length so that there is a smoother transition to the other side of the dependency, which is relatively predictable since it shares high mutual information with the antecedent (Futrell et al., 2019).
Here we do not attempt to adjudicate between the two questions above, nor do we view them in conflict with our proposal. In our opinion, SRA provides a potential explanation for the mechanistic underpinnings of these higher-level processes.
Despite the parallelism, comprehension and production still differ in some critical aspects, exerting modality-specific constraints on SRA due to their idiosyncratic processing nature. For example, production is, in general, more cognitively demanding, in need of higher memory capacity, executive control, and action planning than comprehension (MacDonald, 2013; Nozari & Novick, 2017; Hickok, 2012; Koranda et al., 2020). This additional cognitive demand may exert more pressure to efficiently use the limited memory resources in production than in comprehension, possibly resulting in a stronger effect of SRA. Future work is needed to investigate this possibility.
@@ 7.7 Implications for linguistic typology
Dependency length minimization as a functional universal has been argued to shape the syntactic structure of human language, due to the pressure to efficiently use the limited working memory resources. In the current study, we go beyond the general constraint of limited memory capacity, and further argue that memory resources should be strategically allocated to prioritize novel and unexpected information, a memory efficiency principle that naturally arises from two assumptions about working memory. Our results indicate that this strategic resource allocation indeed serves as a functional constraint to shape syntactic structures, in the sense that the pressure to minimize dependency length can actually be relaxed when the antecedent of a syntactic dependency is of higher surprisal. Our finding further substantiates the functionalist view as a promising approach to provide explanatory accounts for linguistic universals (Gibson et al., 2019). It also highlights the importance of having increasingly sophisticated characterization of functional constraints, in order to see how far we can go with this functionalist view on the structure of human language.
Despite this goal of having SRA as a universal efficiency principle to explain language structure, an important question is: how universal is SRA cross-linguistically, and how does it interact with other grammatical constraints and language-specific phrase structures?
@@
First of all, one consistent pattern we have observed is the asymmetry between subject and object relations. Specifically, the effect of SRA is generally less reliable for object relations than for subject relations. One possible explanation is that object relations are subject to stronger grammatical constraints, with greater pressure to position the head and its dependent closer to each other. In support of this possibility, Keenan & Comrie (1977) identify an Accessibility Hierarchy as a linguistic universal, where noun phrases in the subject position are more readily relativized into relative clauses than those in the object position. Such constraints may bind object noun phrases and verbs more tightly, reducing the influence of SRA. Moreover, in many languages, grammatical agreement is common in subject–verb dependencies but absent in object-verb dependencies. Therefore, establishing subject relations may require more active grammatical computation, resulting in increased memory demand in processing and a stronger need for more strategic and efficient use of working memory resources.
@@
Another notable pattern in Study 1 is that, compared to head-initial languages, most of the head-final languages in our analysis do not exhibit a reliable antecedent surprisal effect on dependency length. We speculate that this may be related to the tendency for argument dropping in head-final languages with SOV word order. From the perspective of dependency locality, SOV word order is associated with longer dependency lengths compared to SVO, which should theoretically impose higher memory costs and make it a less efficient structure. Despite this inefficiency, typologically, SOV is a word order commonly attested (Hammarström, 2016). As an explanation for this paradox, some studies find that arguments in SOV languages are often dropped, reducing the overall dependency length in actual language use (Levshina, 2025; Ueno & Polinsky, 2009). It is possible that by allowing speakers to drop arguments, the efficiency of SOV structure may already be significantly improved, obviating the need for SRA as another efficiency strategy.
@@
7.8 Surprisal effect as efficient coding: An adaptionist view